OpenStudy (anonymous):

Can someone help and explain to me how to do this problem, and work me through the steps??? -2x+4= 2(4x-3) -3 (-8+4x A.x=7 B.x=2 C.x=-7 D.x=3

OpenStudy (austinl):

Ok, we have this: \(-2x+4= 2(4x-3) -3 (-8+4x)\) We need to distribute the constants through the parenthesis. \(-2x+4=8x-6+24-12x\) Combine like terms. \(-2x+4=-4x+18\) Then we need to get variables on one side and constants on the other. \(2x=14\) Then we divide as per usual. \(x=7\)

OpenStudy (anonymous):

ok thank you so much you are very helpful!!!!!! :)